Session: May 2000 (session 2)
Wavelength: 6 cm


To give feedback on this experiment, click here

Experiment Feedback

\(Other\) (Zacchiroli at Sun Jun 4 13:55:26 METDST 2000) Success

Hartebeesthoek (Jonathan Quick at Sun Jun 4 14:15:33 METDST 2000) Success
No problems

Shanghai (Huang Xinyong at Sun Jun 4 14:19:02 METDST 2000) Success

Urumqi (Liu X. at Mon Jun 5 07:46:08 METDST 2000) Minor failures
Maybe read head problem with large parity error. The absulute offset of read is set to -94 microns in head.ctl. The cables of read module was exchanged.

Onsala (A. Polatidis at Mon Jun 5 11:03:36 METDST 2000) Success

Jodrell_Bank (Alastair Gunn at Mon Jun 5 12:27:33 METDST 2000) Success
No known problems.

Cambridge (Alastair Gunn at Mon Jun 5 12:27:58 METDST 2000) Minor failures
Recorder communication problems at start of experiment. Finally working for start of scan at 09:16 UT. No other known problems.

Torun (K.Borkowski at Mon Jun 5 19:23:14 METDST 2000) Minor failures
At 9:22 UT the first of a series of power failures occured. Thus no useful data for almost 2 hours.

Effelsberg (A. Kraus at Tue Jun 6 11:12:48 METDST 2000) Minor failures
The phasecal was not working.

Noto (C.Trigilio at Tue Jun 6 12:45:47 METDST 2000) Success
Midob and postob procedures modified to get continuous tsys measurements.

Westerbork (Tony Foley at Wed Jun 7 13:37:51 METDST 2000) Success
Tied array with 14 telescopes